OpenStudy (anonymous):

If someone could check my answer for 2/7+3/9+5/6=? I came up with 1 71/378 is that right?

OpenStudy (anonymous):

hey you can actually check it in your calculator if you want to

OpenStudy (anonymous):

that would be cool

OpenStudy (anonymous):

i dont have a scientific calculator

OpenStudy (sandra):

hmmm, ok, well one common denominator is definitely 7*9*6 = 378

OpenStudy (sandra):

so it would be: (2*9*6) + (3*7*6) + (5*7*9) all over 378

OpenStudy (anonymous):

add 2/7 and 3/9 first, then reduce to lowest terms. Add that to 5/6 and you should get a denominator of 126

OpenStudy (anonymous):

oh yeah forgot about reducing, without it reduced though it came out to be549/ 378, made to a proper fraction of 1 71/378

OpenStudy (anonymous):

2/7 + 3/9 = 39/63 divided by 3 comes to 13/21. so you can now add 13/21 + 5/6

OpenStudy (sandra):

right, akospiya is telling you a good way to go about it

OpenStudy (sandra):

and btw - both 549 and 278 are divisible by 3

OpenStudy (sandra):

so you could reduce before making the mixed fraction

OpenStudy (anonymous):

appreciate the help, one step closer to remembering this stuff

OpenStudy (radar):

Did you get an answer of 1 19/42?

OpenStudy (anonymous):

Yup